# Correction to Hawking radiation in non-singular gravitational collapse

### v2

@@ I Introduction
The first two assumptions constitute the essence of quantum field theory on curved spacetime. Before Hawking, it was already known through the work of Parker [2] and Imamura [3] that curved spacetimes can induce spontaneous emission of quanta of a quantum field. Indeed, it was understood from the earliest days of quantum field theory that particle production by spontaneous emission can occur in quantum fields in a strong electric field – the so-called Schwinger effect [4, 5, 6]. Thus, it is not surprising that a gravitational background, including that describing a collapsing body, should induce particle production in a quantum field. What turned out to be surprising was the fact that, at late times, the spectrum of emitted particles is thermal, contradicting Hawking's own initial expectation of there being an initial burst of radiation which will eventually die out [7, Chapter 2]. This fact owes its origin to the third assumption referred to above. According to general relativity, given generic initial data, a reasonable matter configuration undergoing gravitational collapse in spherical symmetry will eventually form a Schwarzschild black hole [8, 9, 10]. With this assumption, one can show that the late-time limit of the spectrum of particles emitted by the collapsing matter is thermal [11]. (There are many complementary ways to understand the origin of thermality in radiation from a black hole. The explanation here is tailored for the purposes of this paper; for alternative perspectives, see, e.g., Refs. [12, 13].)
@@
In view of the foregoing discussion, one may ask the following question: to what extent does the spectrum of radiation change if the picture of classical gravitational outlined above is modified? In this connection one must first ask as to why one might expect the classical picture to change. Such an expectation arises from considerations concerning quantum gravity. One of the aims of formulating a satisfactory theory of quantum gravity is to understand the behavior of nature in domains where classical theories of gravity fail. One such domain is the very late stage of gravitational collapse, for then the matter inside the (trapping) horizon approaches the central singularity where general relativity breaks down. Just as the development of a fully consistent quantum theory of electrodynamics sounded a death knell for the infinities plaguing the classical theory of electrodynamics, one expects a fully consistent theory of quantum of gravity to cure the singularities of classical general relativity, and thereby change the conventional picture of such processes as gravitational collapse and the big bang. In fact, investigations into these processes from a variety of quantum theoretical perspectives have yielded promising results. For instance, quantum mechanical evolution of a collapsing thin shell of matter [14, 15, 16, 17] leads to the resolution of the central singularity. More complicated matter profiles require a field-theoretical analysis, which can be accomplished within the framework of various candidate theories of quantum gravity, such as asymptotically free quantum gravity [18, 19, 20] and loop quantum gravity (LQG) [21, 22, 23, 24, 25, 26, 27, 28, 29, and references therein]. The common lesson which emerges from these studies is that gravity becomes repulsive at the Planck scale, portending the demise of cosmological and black-hole singularities, replacing them instead with a bounce in the geometry at a critical radius. Of particular interest are the detailed studies of quantum gravitational collapse from the perspective of asymptotic freedom [18, 19, 20] and of LQG [21, 30, 24]. Despite the difference in approach, both sets of investigation lead to a similar picture of quantum-corrected gravitational collapse (Fig. 1): matter undergoing collapse falls into a transient trapping horizon, thus forming a black hole, bounces outward upon reaching a minimum radius proportional to the Planck length, and eventually exits the trapping horizon, thus bringing the black hole's life to an end  [^1] Refs. [21, 30, 24] also predict a shockwave propagating as a discontinuity in the spacetime geometry. We shall ignore this complication in our computations; however, in the concluding section, we speculate on the potential implications of incorporating the shock-wave discontinuity into the analysis of black hole evaporation. – all of this occurring in a single asymptotic region. This picture of gravitational collapse is radically different from the classical one, and it is the aim of this paper to study how the former affects the spectrum of radiation from the collapsing body.
@@
The paper is organized as follows. In Section II, we review the class of metrics which describe and the tools which enable us to study dynamical black holes, defined as the regions bounded by future trapping horizons – the discussion will be restricted to spherical symmetry. In Section III, the proper-time formalism of Feynman [36, Appendix A] that recasts scalar quantum field theory in terms of the quantum mechanics of a relativistic particle is reviewed, since it forms a precursor to understanding the use of tunneling methods in studying Hawking radiation. These methods are reviewed in Section IV, and used to calculate the probability of spontaneous emission of particles by the trapped region in Fig. 1. Finally, some conclusions from our study are noted in the concluding section.
@@ II Black-hole dynamics
The precise form of $M$ will determine the precise details of how matter behaves (the blue line in the Fig. 1). Examples of mass functions with the bouncing behavior shown in the figure can be found in Refs. [40, 30, 21]. Our calculations will only depend on the generic bouncing behavior of $M(t,r)$, regardless of its precise form.
First of all, following [31, 32], let us study the trapped region in detail, for that is crucial in calculating and understanding the nature of spontaneous emission of particles by the black hole. We start by defining outgoing and ingoing null vectors along the two null directions normal to the 2-spheres foliating the spacetime. To this end, note that along a radial null ray
$\mathrm{d}s^{2}=0$ $=-\mathrm{d}t^{2}+(\mathrm{d}r+N^{r}\mathrm{d}t)^{2}\Rightarrow\frac{\mathrm{d}r}{\mathrm{d}t}=\pm 1-N^{r}\leavevmode\nobreak\ ,$
@@
$n^{a}_{+}=\frac{1}{\sqrt{2}}(1,1-N^{r},0,0)\leavevmode\nobreak\ ,\quad n^{a}_{-}=\frac{1}{\sqrt{2}}(1,-1-N^{r},0,0)\leavevmode\nobreak\ .$ (II.3)
@@
$\theta_{\pm}=q_{ab}\nabla^{a}n^{b}_{\pm}\leavevmode\nobreak\ ,$
@@
$\theta_{\pm}=\pm(1\mp N^{r})$ (II.4)
The trapped region shown in Fig. 1 is bounded by the future trapping horizon, which is defined as the region where the expansion of outgoing null geodesics vanishes, $\theta_{+}=0$, and the expansion of ingoing null geodesics is negative, $\theta_{-}<0$; these two conditions are met by the hypersurface $r=r_{H}(t)$ given by
$r_{H}(t)=2M(t,r_{H}(t))\equiv 2M_{H}\leavevmode\nobreak\ .$ (II.5)
@@
$\frac{{\rm d}r_{H}}{{\rm d}t}=2\dot{M}_{H}+2M^{\prime}_{H}\frac{{\rm d}r_{H}}{{\rm d}t}\leavevmode\nobreak\ ,\qquad\dot{f}\equiv\frac{\partial f}{\partial t}\leavevmode\nobreak\ ,\quad f^{\prime}\equiv\frac{\partial f}{\partial r}\leavevmode\nobreak\ .$ (II.6)
@@
$\frac{{\rm d}r_{H}}{{\rm d}t}=\frac{2\dot{M}_{H}}{1-2M^{\prime}_{H}}\leavevmode\nobreak\ .$ (II.7)
@@
${\rm d}s^{2}|_{H}$ $=\frac{{\rm d}r_{H}}{{\rm d}t}\left(\frac{{\rm d}r_{H}}{{\rm d}t}+2\right){\rm d}t^{2}\leavevmode\nobreak\ +r_{H}{\rm d}\Omega^{2}$ (II.8)
$=\frac{4\dot{M}_{H}}{(1-2M^{\prime}_{H})^{2}}\left[1+(\dot{M}_{H}-2M^{\prime}_{H})\right]{\rm d}t^{2}+r_{H}{\rm d}\Omega^{2}\leavevmode\nobreak\ .$ (II.9)
Outside the collapsing star, and also at its boundary $r=r_{b}(t)$, $M(t,r)=M_{\text{ADM}}$, and so $\dot{M}_{H}=M^{\prime}_{H}=0$, which entails that ${\rm d}s^{2}|_{H}=0$ for constant angular coordinates; this explains why $H_{O}$, the outer part of trapping horizon, which is in the vacuum region, is shown to be null in Fig. 1; in technical terms, it is a non-expanding horizon [41]. The rest of the trapping horizon will have both spacelike and timelike regions, depending on the mass function.
Eqn. (II.9) can be written in a much more illuminating form. In spherical symmetry, there always exists a vector field $K^{a}$, called the Kodama vector field [42, 32], such that $\nabla^{a}(G_{ab}K^{b})=0$. In PG coordinates, it is given by
$K^{a}=(1,0,0,0)\leavevmode\nobreak\ ,$ (II.10)
@@
$K^{a}K_{a}=-1+N^{r}(t,r)^{2}=\theta_{+}\theta_{-}\leavevmode\nobreak\ .$ (II.11)
@@
$\kappa_{H}=\frac{1}{2}g^{AB}\nabla_{A}\nabla_{B}r\leavevmode\nobreak\ |_{H},$ (II.13)
@@
$\kappa_{H}=\frac{1}{4M_{H}}\left[1+(\dot{M}_{H}-2M^{\prime}_{H})\right]\leavevmode\nobreak\ ,$ (II.14)
and so, eqn. (II.9) becomes
${\rm d}s^{2}|_{H}=\frac{16M_{H}\dot{M}_{H}}{(1-2M^{\prime}_{H})^{2}}\kappa_{H}\leavevmode\nobreak\ {\rm d}t^{2}+r_{H}{\rm d}\Omega^{2}\leavevmode\nobreak\ .$ (II.15)
@@
$\mathcal{L}_{n_{-}}\theta_{+}|_{H}=-\sqrt{2}\kappa_{H}\leavevmode\nobreak\ ,$ (II.16)

### v3

@@ I Introduction
The first two assumptions constitute the essence of quantum field theory on curved spacetime. Before Hawking, it was already known through the work of Parker [65] and Imamura [54] that curved spacetimes can induce spontaneous emission of quanta of a quantum field. Indeed, it was understood from the earliest days of quantum field theory that particle production by spontaneous emission can occur in quantum fields in a strong electric field – the so-called Schwinger effect [68, 46, 69]. Thus, it is not surprising that a gravitational background, including that describing a collapsing body, should induce particle production in a quantum field. What turned out to be surprising was the fact that, at late times, the spectrum of emitted particles is thermal, contradicting Hawking's own initial expectation of there being an initial burst of radiation which will eventually die out [41, Chapter 2]. This fact owes its origin to the third assumption referred to above. According to general relativity, given generic initial data, a reasonable matter configuration undergoing gravitational collapse in spherical symmetry will eventually form a Schwarzschild black hole [20, 21, 80]. With this assumption, one can show that the late-time limit of the spectrum of particles emitted by the collapsing matter is thermal [32]. (There are many complementary ways to understand the origin of thermality in radiation from a black hole. The explanation here is tailored for the purposes of this paper; for alternative perspectives, see, e.g., Refs. [83, 63].)
@@
In view of the foregoing discussion, one may ask the following question: to what extent does the spectrum of radiation change if the picture of classical gravitational outlined above is modified? In this connection one must first ask as to why one might expect the classical picture to change. Such an expectation arises from considerations concerning quantum gravity. One of the aims of formulating a satisfactory theory of quantum gravity is to understand the behavior of nature in domains where classical theories of gravity fail. One such domain is the very late stage of gravitational collapse, for then the matter inside the (trapping) horizon approaches the central singularity where general relativity breaks down. Just as the development of a fully consistent quantum theory of electrodynamics sounded a death knell for the infinities plaguing the classical theory of electrodynamics, one expects a fully consistent theory of quantum of gravity to cure the singularities of classical general relativity, and thereby change the conventional picture of such processes as gravitational collapse and the big bang. In fact, investigations into these processes from a variety of quantum theoretical perspectives have yielded promising results. For instance, quantum mechanical evolution of a collapsing thin shell of matter [67, 35, 81, 76] leads to the resolution of the central singularity. More complicated matter profiles require a field-theoretical analysis, which can be accomplished within the framework of various candidate theories of quantum gravity, such as asymptotically free quantum gravity [10, 11, 60] and loop quantum gravity (LQG) [52, 53, 82, 22, 14, 8, 36, 37, 34, 33, 1, and references therein]. The common lesson which emerges from these studies is that gravity becomes repulsive at the Planck scale, portending the demise of cosmological and black-hole singularities, replacing them instead with a bounce in the geometry at a critical radius. Of particular interest are the detailed studies of quantum gravitational collapse from the perspective of asymptotic freedom [10, 11, 60] and of LQG [52, 51, 22]. Despite the difference in approach, both sets of investigation lead to a similar picture of quantum-corrected gravitational collapse (Fig. 1): matter undergoing collapse falls into a transient trapping horizon, thus forming a black hole, bounces outward upon reaching a minimum radius proportional to the Planck length, and eventually exits the trapping horizon, thus bringing the black hole's life to an end  [^1] Refs. [52, 51, 22] also predict a shockwave propagating as a discontinuity in the spacetime geometry. We shall ignore this complication in our computations; however, in the concluding section, we speculate on the potential implications of incorporating the shock-wave discontinuity into the analysis of black hole evaporation. – all of this occurring in a single asymptotic region. This picture of gravitational collapse is radically different from the classical one, and it is the aim of this paper to study how the former affects the spectrum of radiation from the collapsing body.
@@
The paper is organized as follows. In Section II, we review the class of metrics which describe and the tools which enable us to study dynamical black holes, defined as the regions bounded by future trapping horizons – the discussion will be restricted to spherical symmetry. In Section III, the proper-time formalism of Feynman [29, Appendix A] that recasts scalar quantum field theory in terms of the quantum mechanics of a relativistic particle is reviewed, since it forms a precursor to understanding the use of tunneling methods in studying Hawking radiation. These methods are reviewed in Section IV, and used to calculate the probability of spontaneous emission of particles by the trapped region in Fig. 1. Finally, some conclusions from our study are noted in the concluding section. Unless otherwise stated, we work in units in which $G=c=1$.
@@ II Black-hole dynamics
The precise form of $M$ will determine the precise details of how matter behaves (the blue line in the Fig. 1). Examples of mass functions with the bouncing behavior shown in the figure can be found in Refs. [47, 51, 52]. Our calculations will only depend on the generic bouncing behavior of $M(t,r)$, regardless of its precise form. Nevertheless, to spell out the kind of assumptions involved in obtaining a quantum-corrected bouncing model, let us consider a specific example, namely quantum-corrected, marginally bound dust collapse in the framework of effective LQG [51].
II.1 Quantum-corrected dust collapse in LQG
In canonical LQG, the classical starting point is a Hamiltonian formulation of general relativity in terms of $\mathfrak{su}(2)$-valued triads and the Ashtekar-Barbero connection, defined on the spatial hypersurfaces of an arbitrary $3+1$ foilation of a four-dimensional spacetime [6]. One then considers the integrals of the triads over two-dimensional surfaces and holonomies of the Ashtekar-Barbero connection around edges embedded in the spatial hypersurfaces. These objects satisfy the so-called holonomy-flux algebra, which is then represented on a Hilbert space possessing a background-independent inner product. The solution to the quantum theory then consists in representing the diffeomorphism and Hamiltonian constraints of general relativity as operators on the background-independent Hilbert space and à la Dirac, finding the null space of these operators [^2] This part of the LQG program is still not fully complete, the fundamental problem being a fully satisfactory quantization of the Hamiltonian constraint (see Ref. [7] for more details).. The semiclassical limit of the theory can be obtained by finding the expectation values of the holonomy, flux and constraint operators and the quantum equations of motion in suitably defined coherent states [73]. In principle, this should give rise to an effective Hamiltonian field theory which incorporates corrections of order $\hbar$ or higher in the effective equations of motion. In practice, however, taking the semiclassical limit in this way is a delicate business, for one has to deal carefully with problems concerning covariance, anomalies and gauge-fixing (see Ref. [15] for an extensive review in the case of quantum cosmology). However, in the case of marginally bound dust collapse with the gauge-fixed metric (II.1), there is a consistent effective theory, which we now briefly describe [^3] For more general frameworks within LQG, see Refs. [33, 86, 1]. [51].
In marginally bound dust collapse, the Einstein equations in PG coordinates reduce to a single equation for the mass function $M(t,r)$, namely
$\partial_{t}M_{c}-\sqrt{\frac{2M_{c}}{r}}\partial_{r}M_{c}=0\penalty 10000\ ,$ (II.3)
where $c$ indicates that we are referring to a classical function (as opposed to a quantum-corrected function in the effective theory). In this gauge-fixed setting, the triad variables of canonical LQG are all gauge-fixed, whereas the Ashtekar-Barbero connection reduces to a single function $b(t,r)$, the component of the connection which gives rise to holonomies in the $\theta$ direction. It is related to the mass function [^4] For simplicity, we set the Barbero-Immirzi parameter to 1. via
$M_{c}(t,r)=\frac{rb^{2}}{2}$ (II.4)
and hence the equation of motion in terms of $b$ is [^5] In writing down the effective equation of motion for $b$, we have chosen the positive root of $M=rb^{2}/2$. Thus, the $b$ variable that we use differs from that in Ref. [51] by a minus sign.
$\partial_{t}b-\frac{1}{2r}\partial_{r}(rb^{2})=0\penalty 10000\ .$ (II.5)
In passing to the quantum-corrected effective theory, this equation is replaced by [51]
$\partial_{t}b-\frac{1}{2r\mu}\partial_{r}\left(r^{3}\sin^{2}\frac{\sqrt{\mu}b}{r}\right)=0\penalty 10000\ ,$ (II.6)
where $\mu$ is a constant proportional to the square of the Planck length (i.e., $\hbar$ in $G=c=1$ units). From this, we can extract an effective mass function
$M(t,r)=\frac{1}{2\mu}r^{3}\sin^{2}\left(\frac{\sqrt{\mu}b}{r}\right)\penalty 10000\ .$ (II.7)
It can be verified that in the limit $\mu\to 0$, this function reduces to the classical mass function $M_{c}$. It is also clear how this $\mu$-dependent quantum correction can tame classical singularity. From the fact that $\sin^{2}(x)\leq 1$ for all $x\in\mathbb{R}$, it follows that
$M\leq\frac{r^{3}}{2\mu}$ (II.8)
Thus, the mass function is bounded from above, unlike the classical case. The same conclusion can be reached with regard to the energy density of dust, which in PG coordinates is given by
$\rho(t,r)=\frac{\partial_{r}M}{4\pi r^{2}}=\frac{1}{8\pi\mu r^{2}}\partial_{r}\left(r^{3}\sin^{2}\frac{\sqrt{\mu}b}{r}\right)\penalty 10000\ .$ (II.9)
This also allows us to write (II.6) as
$\partial_{t}b=4\pi r\rho\penalty 10000\ ,$ (II.10)
which is useful because it establishes that $b$ is a monotonically increasing function of time (since we may assume $\rho\geq 0$ for all physical matter profiles). We can also extract the domain of $b$ by taking the square-root of (II.7),
$\sin(\sqrt{\mu}b/r)=\sqrt{2\mu M/r^{3}}\penalty 10000\ ,$ (II.11)
the positive root chosen in agreement with (II.6). Since $M\geq 0$ for physical matter, such as dust,
$\frac{\sqrt{\mu}b}{r}\in[0,\pi]\penalty 10000\ .$ (II.12)
To understand the qualitative features of the model, it is instructive to look at a specific mass profile, such as that for the Oppenheimer-Snyder model of stellar collapse [51]. In that case, there is a moving boundary $r=L(t)$ of the star inside which the energy density of dust is radially constant and outside which it is zero. Thus, for $r>L(t)$, we have
$\rho=0\Rightarrow M_{e}(t,r)=M_{\text{ADM}}\penalty 10000\ ,$ (II.13)
where the $e$ in the subscript refers to the exterior region ($r>L(t)$). This implies that in the exterior of the star, $b$ is given by the equation
$\sin\left(\frac{\sqrt{\mu}b_{e}}{r}\right)=\sqrt{\frac{2\mu M_{\text{ADM}}}{r^{3}}}\penalty 10000\ ,$ (II.14)
Since the right-hand side is a positive constant in time, to make sure that $b_{e}\in[0,\pi]$, we can have either
$b_{e}=\frac{r}{\sqrt{\mu}}\arcsin\left(\sqrt{\frac{2\mu M_{\text{ADM}}}{r^{3}}}\right)$ (II.15)
or the same expression subtracted from $\pi$. We choose the former. In the interior region $(r<L(t))$, $\partial_{r}\rho=0$, which implies that $\partial_{r}(\sqrt{\mu}b/r)=0$. Thus, (II.6) becomes [51]
$\partial_{t}\left(\frac{\sqrt{\mu}b_{i}}{r}\right)=\frac{3\sin^{2}(\sqrt{\mu}b_{i}/r)}{2\sqrt{\mu}}\penalty 10000\ ,$ (II.16)
where the $i$ in subscript refers to the interior of the star. This equation has the solution
$\sin(\sqrt{\mu}b_{i}/r)=\frac{1}{\sqrt{1+9(t-t_{0})^{2}/4\mu}}\penalty 10000\$(II.17)
for some constant $t_{0}$. To determine $b_{i}$ from this equation, some care is required. The right-hand side of the preceding equation is now time-dependent, and in particular, increases in time for $t<t_{0}$ and decreases in time for $t>t_{0}$. Therefore, since $b$ must always be monotonically increasing in time, $b_{i}\in[0,\pi/2]$ for $t<t_{0}$ and $b_{i}\in[\pi/2,\pi]$ for $t>t_{0}$. That is, we must have
$b_{i}=\begin{cases}\arcsin\left(\frac{1}{\sqrt{1+9(t-t_{0})^{2}/4\mu}}\right)\penalty 10000\ ,&t<t_{0}\\ \pi-\arcsin\left(\frac{1}{\sqrt{1+9(t-t_{0})^{2}/4\mu}}\right)\penalty 10000\ ,&t>t_{0}\end{cases}$ (II.18)
To determine the evolution of the boundary $L(t)$ of the star, one can try to match the external and internal solution. Doing this for $t<t_{0}$ yields
$L(t)^{3}=\frac{M_{\text{ADM}}}{2}\left(4\mu+9(t-t_{0})^{2}\right)$ (II.19)
However, this only holds $t<t_{0}$, for then both $b_{i},b_{e}\in[0,\pi/2]$, whereas for $t>t_{0}$, $b_{e}\in[0,\pi/2]$ while $b_{i}\in[\pi/2,\pi]$, making it impossible to impose continuity on $b$. This is a manifestation of the fact that (II.6) is a hyperbolic conservation law in the variable $B=rb$, and smooth solutions to hyperbolic conservation laws can develop discontinuities (called shock waves in the technical literature) in a finite time [59]. To the find the solution beyond the appearance of the first discontinuity, one has to integrate the conservation law over $S-\epsilon,S+\epsilon]$, where $r=S$ is the location of the shock wave. This leads to an evolution equation for the shock wave, the so-called Rankine-Hugoniot (RH) condition
$\frac{{\rm d}S}{{\rm d}t}=\frac{S(t)^{2}}{2\mu}\frac{M(t,b^{+})-M(t,b^{-})}{b^{+}-b^{-}}\penalty 10000\ ,$ (II.20)
where $b^{\pm}=\lim_{r\to S^{\pm}}b$. Although this equation is necessary to find the correct dynamics in the rebound phase, for simplicity, whenever using the Oppenheimer-Snyder model, we will assume $L(t)$ to have the same form in the post-bounce phase ($t>t_{0}$) as it has in the pre-bounce phase ($t<t_{0}$), i.e., eqn. (II.19). This simplifying assumption will change the black-hole life time [51] from what it should turn out to be under (II.20), but this equation is just one possible instance of the RH condition. In general, there are many inequivalent ways of integrating a conservation law, and each leads to a different RH condition, which in turn leads to a different lifetime for the black hole [27]. In the concluding section, we shall outline an argument suggesting that Hawking radiation from the collapsing body might disperse the discontinuity in the mass function and hence one may be able to dispense with the need of using the RH condition altogether.
Notice that (II.19) entails that the falling matter does not fall to $r=0$, as in the classical case, but reaches the minimum radius $L_{\text{min}}$, and then rebounds and starts falling outwards. This is also clear from the behavior of the mass function and energy density inside the star. From (II.17), we find that
$M_{i}(t,r)=\frac{2r^{3}}{4\mu+9(t-t_{0})^{2}}\penalty 10000\$(II.21)
and hence
$\rho_{i}=\frac{2}{2\pi(4\mu+9(t-t_{0})^{2})}\penalty 10000\ .$ (II.22)
The bouncing nature of the geometry is manifest. For fixed $r$, say $r=r^{*}$, as $t\to t_{0}^{+}$, $M(t,r)$ increases, i.e., more and more mass piles up in the region $r\leq r^{*}$. When $t=t_{0}$, the mass function attains its maximum value $r^{*3}/2\mu$. And for $t>t_{0}$, the mass function decreases with time, indicating that the dust accumulated behind $r^{*}$ has now started to move out, i.e. the star has bounced. It is also worth emphasizing that the bounce is of quantum origin: it is the constant $\mu$, proportional to $\hbar$, which ensures that the energy density in the interior undergoes a bounce at $t=t_{0}$ rather than diverging as $t\to t_{0}^{+}$, as in the classical theory [^6] In Ref. [51], which we have closely followed in this subsection, the shift function $N^{r}$ is different in the effective theory from what we have assumed. We have introduced a quantum correction in the equation of motion for $b$, which modifies the mass function $M$, which in turn modifies the shift function through the classical relation $N^{r}=\sqrt{2M/r}$. In Ref. [51], however, the quantum correction to $N^{r}$ is introduced independently of the correction to $b$, and this renders the relation $N^{r}=\sqrt{2M/r}$ in their framework. We have chosen to avoid this path because (1) the effective theory allows it (i.e., it is consistent) and (2) the relation $N^{r}=\sqrt{2M/r}$ leads to a very simple equation for the location of the horizon, as shown in the next subsection..
II.2 Dynamical horizons
Let us now return to a generic mass function $M(t,r)$, and ask how such features as the location of trapped regions and the black-hole boundary can be studied in the context of a general time-dependent spacetime that describes gravitational collapse. To this end, we shall follow the ideas first laid out in Refs. [43, 44]. Let us begin by observing that due to spherical symmetry, every $(t,r)$ corresponds to a 2-sphere of symmetry, and every such sphere has two independent – one ingoing and one outgoing – null vectors orthogonal to the sphere. To write down these null vectors, note that along a radial null ray
$\mathrm{d}s^{2}=0$ $=-\mathrm{d}t^{2}+(\mathrm{d}r+N^{r}\mathrm{d}t)^{2}\Rightarrow\frac{\mathrm{d}r}{\mathrm{d}t}=\pm 1-N^{r}\penalty 10000\ ,$
@@ II.2 Dynamical horizons
$n^{a}_{+}=\frac{1}{\sqrt{2}}(1,1-N^{r},0,0)\penalty 10000\ ,\quad n^{a}_{-}=\frac{1}{\sqrt{2}}(1,-1-N^{r},0,0)\penalty 10000\ .$ (II.23)
@@
$\theta_{\pm}=q_{ab}\nabla^{a}n^{b}_{\pm}\penalty 10000\ ,$
@@
$\theta_{\pm}=\pm(1\mp N^{r})$ (II.24)
The trapped region shown in Fig. 1 is bounded by the future trapping horizon, which is defined as the region where the expansion of outgoing null geodesics vanishes, $\theta_{+}=0$, and the expansion of ingoing null geodesics is negative, $\theta_{-}<0$; these two conditions are met if $N^{r}=1$, which is true at the hypersurface $r=r_{H}(t)$ given by
$r_{H}(t)=2M(t,r_{H}(t))\equiv 2M_{H}\penalty 10000\ .$ (II.25)
@@
$\frac{{\rm d}r_{H}}{{\rm d}t}=2\dot{M}_{H}+2M^{\prime}_{H}\frac{{\rm d}r_{H}}{{\rm d}t}\penalty 10000\ ,\qquad\dot{f}\equiv\frac{\partial f}{\partial t}\penalty 10000\ ,\quad f^{\prime}\equiv\frac{\partial f}{\partial r}\penalty 10000\ .$ (II.26)
@@
$\frac{{\rm d}r_{H}}{{\rm d}t}=\frac{2\dot{M}_{H}}{1-2M^{\prime}_{H}}\penalty 10000\ .$ (II.27)
@@
${\rm d}s^{2}|_{H}$ $=\frac{{\rm d}r_{H}}{{\rm d}t}\left(\frac{{\rm d}r_{H}}{{\rm d}t}+2\right){\rm d}t^{2}\penalty 10000\ +r_{H}{\rm d}\Omega^{2}$ (II.28)
$=\frac{4\dot{M}_{H}}{(1-2M^{\prime}_{H})^{2}}\left[1+(\dot{M}_{H}-2M^{\prime}_{H})\right]{\rm d}t^{2}+r_{H}{\rm d}\Omega^{2}\penalty 10000\ .$ (II.29)
Outside the collapsing star, and also at its boundary $r=r_{b}(t)$, $M(t,r)=M_{\text{ADM}}$, and so $\dot{M}_{H}=M^{\prime}_{H}=0$, which entails that ${\rm d}s^{2}|_{H}=0$ for constant angular coordinates; this explains why $H_{O}$, the outer part of trapping horizon, which is in the vacuum region, is shown to be null in Fig. 1; in technical terms, it is a non-expanding horizon [9]. The rest of the trapping horizon will have both spacelike and timelike regions, depending on the mass function. Since $\dot{M}_{H_{I}}<0$ ($\dot{M}_{H_{I}}>0$) in the collapse (rebound) phase, we should generally expect the inner part $H_{I}$ of the horizon to be timelike before the bounce and spacelike after the bounce.
Eqn. (II.29) can be written in a much more illuminating form. In spherical symmetry, there always exists a vector field $K^{a}$, called the Kodama vector field [56, 44], such that $\nabla^{a}(G_{ab}K^{b})=0$. In PG coordinates, it is given by
$K^{a}=(1,0,0,0)\penalty 10000\ ,$ (II.30)
@@
$K^{a}K_{a}=-1+N^{r}(t,r)^{2}=\theta_{+}\theta_{-}\penalty 10000\ .$ (II.31)
@@
$\kappa_{H}=\frac{1}{2}g^{AB}\nabla_{A}\nabla_{B}r\penalty 10000\ |_{H},$ (II.33)
@@
$\kappa_{H}=\frac{1}{4M_{H}}\left[1+(\dot{M}_{H}-2M^{\prime}_{H})\right]\penalty 10000\ ,$ (II.34)
and so, eqn. (II.29) becomes
${\rm d}s^{2}|_{H}=\frac{16M_{H}\dot{M}_{H}}{(1-2M^{\prime}_{H})^{2}}\kappa_{H}\penalty 10000\ {\rm d}t^{2}+r_{H}{\rm d}\Omega^{2}\penalty 10000\ .$ (II.35)
@@
$\mathcal{L}_{n_{-}}\theta_{+}|_{H}=-\sqrt{2}\kappa_{H}\penalty 10000\ ,$ (II.36)